Make or Buy Decision:

Zee-Drive Ltd. is a computer manufacturer. One of the items they make is monitors. Zee-Drive has the opportunity to purchase 20,000 monitors from an outside supplier for $197 per unit. One of the company's cost-accounting interns prepared the following schedule of Zee-Drive's cost to produce 20,000 monitors:

Total cost of producing 20,000 monitors

Unit cost
Direct materials $ 2,320,000 $ 116
Direct labor 1,320,000 66
Variable factory overhead 640,000 32
Fixed manufacturing overhead 500,000 25
Fixed non-manufacturing overhead 840,000 42
$ 5,620,000 $ 281

You are asked to look over the intern's estimate before the information is shared with members of management who will decide to continue to make the monitors or buy them. The company's controller believes that the estimate may be incorrect because it includes costs that are not relevant. If Zee-Drive buys the monitors, the direct labor force currently employed in producing the monitors will be terminated and there would be no termination costs incurred. There are no materials on hand and no commitments to suppliers to purchase materials, so all materials would need to be purchased to make the monitors. Variable overheads are avoidable if monitors are bought. Fixed manufacturing overhead costs would be reduced by $40,400, but non-manufacturing costs would remain the same if monitors are bought.

Fill in the differential analysis.

Make or Buy Decisions
Differential Analysis Report
Purchase price of 20,000 monitors $
Differential cost to make:
Direct materials $
Direct labor
Overhead
Differential income (loss) from making monitors $

Feedback

Enter only the differential relevant costs in the appropriate space and calculate differential income or loss. The challenge is in determining differential overhead.

Keep or Replace Machine:

Skiles Coporation is a manufacturer of classic rocking chairs. The company has been using a particular sanding and finishing machine for over 10 years and believes that it may be time to replace the machine. The company is trying to decide whether replacing the old machine is a wise economic decision. The company's controller pulled together the following information on the old machine and the new possible replacement machine.

Old Machine:
Original cost $452,700
Current accumulated depreciation 301,100
Estimated annual variable manufacturing costs for machine 73,250
Estimated selling price of machine 198,300
Estimated remaining useful life (in years) 6
New Machine:
Purchase cost $814,600
Estimated annual variable manufacturing costs for machine 49,350
Estimated residual value 0
Estimated useful life (in years) 6

Select the relevant or irrelevant information below:

Annual variable costs of old machine Relevant
Selling price of old machine Relevant
Matching lives Relevant
Purchase price of new machine Relevant
Accumulated depreciation of old machine Irrelevant

Fill in the differential analysis.

Replace or Keep Decision
Differential Analysis Report
Cost of replacing old machine:
Annual differential decrease in cost $
x number of years
Total differential decrease in cost
Proceeds from sale of present machine $
Cost of new machine
Net differential (increase)/decrease in cost, six year total $
Answer #1

Part 1 Zee-Drive Ltd.

Make or Buy Decisions
Differential Analysis Report

Purchase price of 20,000 monitors

$3940000

Differential cost to make:

Direct materials

$2320000

Direct labor

1320000

Overhead

Variable = 640000

Fixed = 40400

680400

3926400

Differential income (loss) from making monitors

$(380400)

Purchase price = 20000*197 = 3638000

Direct material= 20000*116 = $2320000

Direct labor = 20000*66 = 1320000

Variable factory overhead = 20000*32 = 640000

Part B

Skiles Coporation

Annual variable costs of old machine

· Relevant

Selling price of old machine

· Relevant

Matching lives

Relevant

Purchase price of new machine

Relevant

Accumulated depreciation of old machine

· Irrelevant

Replace or Keep Decision
Differential Analysis Report

Cost of replacing old machine:

Annual differential decrease in cost

$23900 (73250-49350)

x number of years

6

Total differential decrease in cost

$143400

Proceeds from sale of present machine

198300

Cost of new machine

814600

Net differential (increase)/decrease in cost, six year total

$472900

Net differential decrease= Cost of new machine- Total differential decrease in cost- Proceeds from sale of present machine